# Beyond Isolation: Multi-Agent Synergy for Improving Knowledge Graph Construction

### v2

@@ Abstract.
Knowledge graph construction (KGC) is a multifaceted undertaking involving the extraction of entities, relations, and events. Traditionally, large language models (LLMs) have been viewed as solitary task-solving agents in this complex landscape. However, this paper challenges this paradigm by introducing a novel framework, CooperKGC. Departing from the conventional approach, CooperKGC establishes a collaborative processing network, assembling a KGC collaboration team capable of concurrently addressing entity, relation, and event extraction tasks. Our experiments unequivocally demonstrate that fostering collaboration and information interaction among diverse agents within CooperKGC yields superior results compared to individual cognitive processes operating in isolation. Importantly, our findings reveal that the collaboration facilitated by CooperKGC enhances knowledge selection, correction, and aggregation capabilities across multiple rounds of interactions [^1] All code are available at https://github.com/hongbinye/CooperKGC..
Knowledge Graph Construction, Information Extraction, Multi-agent Collaboration
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Information extractionccs: Computing methodologies Cooperation and coordination
1. Introduction
@@
In the era of information abundance, constructing comprehensive knowledge graphs (Ye et al., 2022; Mondal et al., 2021; Vakaj et al., 2023) has emerged as a pivotal task, facilitating the organization and retrieval of structured information essential for various applications (Suárez and Hogan, 2023; Atif et al., 2023; Jiang and Usbeck, 2022). Prior research (Lu et al., 2022; Paolini et al., 2021; Lu et al., 2021) in the domain has predominantly focused on fine-tuning task-specific models through extensive datasets, wielding the power of supervised learning. The advent of LLMs, such as GPT-3 (Brown et al., 2020) and ChatGLM (Du et al., 2022), has revolutionized natural language processing by showcasing unparalleled proficiency in understanding and generating human-like text. However, the application of these models to KGC remains an intricate challenge, as this task necessitates not only language understanding but also precise extraction of elements within the confines of predefined schemas. Recent investigations (Li et al., 2023c; Zhu et al., 2023; Gao et al., 2023) have revealed that, despite their triumphs in various natural language tasks, LLMs still grapple with performance shortcomings when confronted with the intricate demands of KGC tasks. Notably, Wei et al. (2023) reveals that the raw textual data utilized to train large language models may lack task-specific schemas, resulting in a weakened semantic grasp and structural analysis of the underlying schema. Therefore we contends that a shift from traditional parameter-based paradigms to a more nuanced approach, like Chain-of-Thought (CoT) (Wei et al., 2022; Yu et al., 2023), can address the challenges posed by multi-step inference problems inherent in KGC. By decomposing the information extraction task into distinct stages, it not only reduces search space but also computational complexity, thereby offering a promising avenue for improving KGC.
Embracing the profound insights from the Society of Mind (SOM) (Minsky, 1988), which conceptualizes the mind as a complex system emerging from the interactions of simple components, our research explores the transformative potential of LLM-based agents in multi-agent systems. These agents (Xi et al., 2023) exhibit cooperative behaviors in diverse scenarios, spanning role-playing environments (Li et al., 2023a), the formation of interactive artificial societies (Park et al., 2023), and task-oriented social simulations (Li et al., 2023d). Taking inspiration from pioneering work of Liang et al. (2023), we employ the multi-agent debate framework for collaborative self-reflection on challenging tasks. Collaboration is defined as an iterative refinement process, wherein each round generates a new answer based on prior answers and self-reflection. This iterative feedback fosters continuous improvement, making our collaborative approach adept at tackling problems that elude single-agent solutions. Through this synergistic collaboration, the agent network transcends individual capabilities, paving the way for KGC.
Figure 1. A knowledge graph construction team which consists of experts proficient in various tasks. In contrast to the previous stage decomposition methods, we construct it as a multi-intelligence collaborative process, which facilitates interaction based on collaborative mindset while introducing other perspectives.
@@
Specifically, our dedicated team of agents comprises experts proficient in various tasks, including named entity recognition, relation extraction, and event extraction. Figure 1 illustrates a scenario where given a sentence ”Born in Rosario, Argentina in 1987 Lionel Messi is widely held to be one of the greatest football players of his generation.” Despite the event extraction agent successfully identifies the ”Life:Be-Born” event triggered by the word ”Born” and recognizes associated arguments ”Lionel Messi” and ”Rosario” with roles ”Person” and ”Place,” respectively, it misses the crucial meta role ”Time” for the argument ”1987.” Notably, the named entity recognition agent identifies ”1987” as a time entity, while the relation extraction agent establishes the triplet ”Lionel Messi, per:date_of_birth,1987.” Through collaborative interactions within the team, the event extraction agent adeptly supplements the missing event arguments. This collaborative approach, inspired by the harmonious interaction (Singh, 2003) advocated by the SoM framework and akin to the dynamics of individuals (Woolley et al., 2010) with different specialties in human society (Sap et al., 2022), demonstrates the advantages of multi-agent collaboration (Miller, 2010; Wuchty et al., 2007). In our approach CooperKGC , we construct a collaborative team of agents, each specializing in distinct tasks to simulate the nuanced teamwork prevalent in human society. The integration of open interaction, expertise refinement, and adaptability to others’ opinions mirrors the foundations of a cohesive society. Our exploration into diverse collaboration strategies reveals key insights: (1) Inclusion of agents with varied expertise enhances collaboration outcomes. (2) While model hallucinations (Ye et al., 2023) may arise, effective communication among team members mitigates these drawbacks. (3) Substantial team collaboration enhances extraction results on target tasks; however, an intriguing observation emerges that increasing cooperation rounds doesn’t invariably yield superior results. In our collaborative mechanism, balancing interaction frequency ensures the expert agent’s beliefs remain undisturbed by excessive external authoritative information, aligning with fundamental theories of sociology (Tuckman, 1965; Goffman et al., 2002; Durkheim, 2018). In summary, our work contributes the following:
  • We introduce a pioneering method, cooperKGC, that capitalizes on the inherent reasoning capabilities of LLMs and orchestrates a multi-agent collaborative process to elevate the precision of existing models. Our empirical findings underscore the efficacy of this approach.
  • We delve into the mechanics of collaboration within multi-agent teams. Our objective is to discern the nuances of team intelligence exhibited by LLMs possessing diverse expertise when engaged in collaborative endeavors. Enriching our study, we draw upon sociological theory to contextualize the behavioral and decision-making tendencies of expert agents within multi-turn conversational interactions.
  • Intriguingly, our simulations, mirroring the subtle collaborative capabilities inherent in human society through LLM agents, unveil the significance of configuring adept team members and adopting efficient collaboration strategies for multi-agent systems. This revelation hints at a future where artificial intelligence becomes inherently collaboration-aware, marking a pivotal stride in the evolution of AI.
2. Related work
@@
AutoKG (Zhu et al., 2023) [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
ChatIE (Wei et al., 2023) [Uncaptioned image] (2-agents) [Uncaptioned image] [Uncaptioned image] (2-rounds) [Uncaptioned image] [Uncaptioned image]
GPT-NER (Wang et al., 2023b) [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] (¿3-rounds) [Uncaptioned image] [Uncaptioned image]
GPT-RE (Wan et al., 2023) [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
CoT-ER (Ma et al., 2023) [Uncaptioned image] (3-agents) [Uncaptioned image] [Uncaptioned image] (3-rounds) [Uncaptioned image] [Uncaptioned image]
LM vs LM (Cohen et al., 2023) [Uncaptioned image] (2-agents) [Uncaptioned image] [Uncaptioned image] (3-stages) [Uncaptioned image] [Uncaptioned image]
Multiagent Debate (Du et al., 2023) [Uncaptioned image] (2-agents) [Uncaptioned image] [Uncaptioned image] (3-stages) [Uncaptioned image] [Uncaptioned image]
MAD (Liang et al., 2023) [Uncaptioned image] (3-agents) [Uncaptioned image] [Uncaptioned image] (3-stages) [Uncaptioned image] [Uncaptioned image]
PRD (Li et al., 2023b) [Uncaptioned image] (2-agents) [Uncaptioned image] [Uncaptioned image] (3-stages) [Uncaptioned image] [Uncaptioned image]
SPP (Wang et al., 2023a) [Uncaptioned image] (¿3-agents) [Uncaptioned image] [Uncaptioned image] (4-stages) [Uncaptioned image] [Uncaptioned image]
@@
Table 1. Comparison with previous methods. The upper half represents LLM-based KGC method, while the lower half showcases the emerging multi-agent approach applied to other tasks. Our CooperKGC pioneers a multi-expert collaborative processing method explicitly tailored for KGC tasks. Notable components include the ”¿3-rounds” concept, indicating iterative interactions beyond three rounds, the ”3-stage” process involving statement, interactive debate, and summary iterations, and the more advanced ”4-stage” introducing dynamic allocation of expert agents.
2.1. LLM-based Knowledge Graph Construction
@@
Knowledge graph construction has evolved significantly from its early dependence on supervised learning methods. Recent years have witnessed a surge of interest in leveraging the remarkable advancements achieved by LLMs within the realm of KG. Notably, Zhu et al. (2023) delves into the application of LLMs in KG construction and reasoning tasks, uncovering their commendable prowess in tasks such as reasoning and question answering. Building on this foundation, Zhang et al. (2023) extends the exploration by integrating KG structural information into LLMs, employing self-supervised structural embedding pre-training to imbue models with structural awareness. Wei et al. (2023), in a novel perspective, posits that LLMs inherently possess the capability for unified zero-shot information extraction in interactive settings. This belief led to the proposal of a two-stage framework that transforms information extraction into a multi-turn question and answer architecture. Furthermore, Pang et al. (2023) pinpoints the unspecified task description as a key factor hindering the performance of contextual information extraction. To address this, a guided learning framework is introduced to enhance the extraction model’s alignment with specified guidelines. In the broader landscape of NLP, diverse approaches have been explored to enhance reasoning in language models, ranging from scratchpad (Nye et al., 2021), verification (Cobbe et al., 2021), CoT (Wei et al., 2022; Kojima et al., 2022; Reynolds and McDonell, 2021), self-reflection (Shinn et al., 2023; Madaan et al., 2023), and fine-tuning (Lewkowycz et al., 2022; Rajani et al., 2019; Zelikman et al., 2022). Techniques such as RLHF (Liu et al., 2022a; Ziegler et al., 2019), knowledge retrieval enhancement (Guu et al., 2020), and likelihood estimation-based training-free methods (Kadavath et al., 2022) have also been employed to bolster the factual consistency. Departing from the conventional isolation of KGC as a singular task, our approach advocates a departure from such isolation by fostering collaboration among a group of expert model agents in a multi-round social environment. This novel perspective encourages models to reflect on their solutions based on the outcomes of tasks performed by other agents, thereby ushering in a paradigm shift towards multi-agent synergy for KGC.
2.2. Interactive Collaboration of Multiple Agents
In the landscape of collaborative efforts among language model agents, a trend has emerged in recent years showcasing the robust performance achieved through the synergy of multiple agents across various tasks, offering a promising avenue to augment the capabilities of individual LLMs. Noteworthy advancements in facilitating collaboration among multiple agents include the development of diverse interaction architectures, often assigning agents to static roles. For instance, Du et al. (2023) employs a two-model agent setup engaging in multiple rounds of debate, resulting in heightened factuality and reasoning capabilities, albeit at the cost of increased computational expenses. Similarly, Cohen et al. (2023) introduces an examiner LLM to validate claims produced by the original LLM, leveraging a distinct division of labor to uncover factual errors. Hao et al. (2023)’s contribution lies in the ChatLLM network, enabling dialogue-based language models to interact, share feedback, and collectively deliberate on problems, fostering diverse perspectives within the system. Furthermore, introducing a judge into the decision-making process, responsible for summarizing debates and rendering final conclusions (Liang et al., 2023; Xiong et al., 2023), demonstrates the potential of guided debates to enhance performance. To Address self-enhancement bias, Li et al. (2023b) proposes a framework incorporating peer review and discussion. While Wang et al. (2023a) explores the role of multiple agents in rounds of self-collaboration, concentrating on refining problem-solving capabilities for a singular complex task. In this study, we inherit previous methods by presenting an interactive architecture tailored for a knowledge graph construction team comprising agents with distinct expert skills. A departure from prior work, our innovation lies in introducing the challenge of completing multiple information extractions, aligning with the principles of feedback compatibility.
3. Preliminaries
In this paper, we apply our approach, cooperKGC, to three representative tasks of knowledge graph construction, namely: Named Entity Recognition (NER), Relation Extraction (RE) and Event Extraction (EE). Notably, since these three tasks are relevant in some way, we leave the exploration of cross-domain collaboration for future work. We provide definitions for these sub-tasks and then explain the basic structure of current models.
3.1. Task Definition
Figure 2. The overview of our CooperKGC.
Named Entity Recognition. Standing as a fundamental and pivotal undertaking in the construction of knowledge graphs, this task involves the identification of entities within texts and their subsequent classification into predefined categories, encompassing names of individuals, geographical locations, proper nouns, organizational designations, etc. A prime example illustrating the significance of NER is evident in the sentence “Messi wins record-extending 8th Ballon d’Or in 2023.”, In this context, the model is tasked with pinpointing the specific entities representing Person (Messi) and Time (2023). The proficiency of NER in discerning such elements is instrumental in extracting meaningful information from textual data and facilitating the organization of KGs.
Relation Extraction. The crucial task of knowledge graph construction involves simultaneously extracting entity mentions and their relations, represented as triples (subject, relation, object), from unstructured texts. Given the input sentences, the desired outputs are relational triples $(e_{h},r,e_{t})$, where $e_{h}$ is the head entity, $r$ is the relation, and $e_{t}$ is the tail entity. For instance, given the sentence “Kylian Mbappé grew up in Bondy, one of the Parisian suburbs with working-class, mostly immigrant residents.”, the model should identify two entities Kylian Mbappé and Bondy, together with their relation person-place-lived, described as triple (Kylian Mbappé, person-place-lived, Bondy).
Event Extraction: The automated extraction of events from unstructured natural language texts, guided by an event schema, constitutes event extraction. To elaborate on the procedure, terms are defined as follows: a trigger word denotes a word or phrase that best characterizes the event, while an event argument represents an entity or attribute associated with the event, such as the place or time. For example, the sentence “Li Shaomin was convicted of espionage and deported.” describes an Convict event triggered by the word ‘convicted’. This event includes two argument roles: the Defendant (Li Shaomin) and the Crime (espionage). The model should be able to identify event triggers, their types, arguments, and their corresponding roles.
3.2. Problem Formulation
In the context of an original text $\mathcal{X}$, the objective is to extract the requisite elements $\mathcal{Y}=\{\mathcal{Y}^{1},...,\mathcal{Y}^{n}\}$, in alignment with the predefined constraints outlined by the schema $\mathcal{S}$. Note that $\mathcal{Y}^{i},i\in t$ represents the information to extract for the i-th type, and $n$ refer to the number of types. For the named entity recognition task, $\mathcal{Y}^{i}$ is in the form of tuples $\mathcal{Y}^{i}=(entity-type,entity-span)$. For the relation extraction task, $\mathcal{Y}^{i}$ is in the form of triples $\mathcal{Y}^{i}=(e_{h},r,e_{t})$, including the head entity, tail entity, and corresponding relation. For the event extraction, $\mathcal{Y}^{i}$ contains the event record in the sentence, which can be represented as $\mathcal{Y}^{i}=\{event-type,event-trigger,argument-span,argument-role\}$. In the following sections, we present the details of team formation and interaction.
4. Methodology
@@
Illustrated in Figure 1, we introduces a collaborative framework CooperKGC, aimed at advancing knowledge graph construction by concurrently extracting component elements such as entities, relations, and events. This framework promotes the interconnection of agents engaged in various tasks through collaborative team efforts (Section 4.1). Building upon this collaborative foundation, we delve into the customization of expert knowledge backgrounds for each agent (Section 4.2). A strategic approach is introduced to multiple rounds of collaborative interaction, aimed at refining individual extraction results by comprehensively incorporating the perspectives of other agents (Section 4.3). Notably, our method could extend beyond the confines of the selected three tasks, offering flexibility through a dynamically formulated team collaboration network tailored to specific task requirements. This adaptive and collaborative approach underscores our commitment to moving ”Beyond Isolation,” fostering multi-agent synergy for an improved and nuanced KGC process.
4.1. Construction of Multi-expert Agent Collaboration Network
In the collaboration process of LLM-agents, our approach advocates a dynamic exchange of text messages through multiple rounds of interactions. Traditional methods often treat expert agents, each equipped with distinct back-ends, as isolated nodes within the collaborative network. These nodes independently contribute to task-solving through separate thinking chains, and a central adjudication node amalgamates and rectifies their responses. However, this conventional solution reveals two critical flaws: (1) The adjudication node, functioning as the central hub, exhibits low fault tolerance and demands substantial reasoning ability to assimilate opinions from nodes spread across diverse collaborative networks. This centralized load imposes elevated requirements on the system; (2) The team heavily relies on the ruling node as the sole consensus mechanism, hindering effective interactions between participants in the KGC task.
@@ 4.1. Construction of Multi-expert Agent Collaboration Network
In response to these limitations, we advocate a decentralized collaborative network communication scheme. Here, each expert agent backend, responsible for handling a specific task, establishes a bidirectional communication channel with any other expert agent backend. Despite the asynchronous nature of message production during practical operations, we adopt rounds as the fundamental unit of interaction to accomplish designated tasks and facilitate replica communication among expert agents. It is noteworthy that, although our approach draws inspiration from the Byzantine Fault Tolerance (Lamport et al., 1982) to form a distributed network, the message records held by each agent node differ. This divergence arises due to the inherent constraints of LLMs related to token length and reasoning performance bottlenecks. In the process of replica communication, we implement message simplification, whereby extraction results complying with schema constraints are distilled. This strategic simplification enhances the efficiency of information exchange and ensures a more seamless and organized collaborative effort in the construction of knowledge graphs. The formalization of the abstract collaboration network comprises three fundamental components:
@@
4.2. Customized Expert Knowledge Background.
@@
Task Definition. The task description $\mathcal{P}_{t}$ can be further decomposed into three components, as exemplified by the RE agent:(1) The first sentence of the task description, "You are an excellent expert in relation extraction." is a constant that tells the LLM that it needs to focus on the relation extraction task; (2) The second sentence defines the output format of the task: "The relation extraction task is given a list of relations in the format (Relation Type: [Head Entity Type, Tail Entity Type]) You need to extract the relation type, the corresponding head and tail entities that may be included in the raw text, or answer NO if they do not exist. Each result is returned as a tuple, e.g. [(head entity 1, relation type 1, tail entity 1), (head entity 2, relation type 2, tail entity 2), ...]". (3) The third sentence points to a specific list of relation types : "The list of constrained relations is: [#Relation 1: [#Head Entity Type 1, #Tail Entity Type 2]...]".
@@
Algorithm 1 The Optimization Process of CooperKGC on an arbitrary input text
Table 2. F1-score results for 3 KGC tasks (NER, RE, EE) on the 8 datasets. The highest scores are in bold, and the underline represents the second-best-performing method. Note that the baseline model results are rerun with consistent settings.

### v3

@@ Abstract
This paper introduces CooperKGC, a novel framework challenging the conventional solitary approach of large language models (LLMs) in knowledge graph construction (KGC). CooperKGC establishes a collaborative processing network, assembling a team capable of concurrently addressing entity, relation, and event extraction tasks. Experimentation demonstrates that fostering collaboration within CooperKGC enhances knowledge selection, correction, and aggregation capabilities across multiple rounds of interactions.
Keywords:
Knowledge graph construction Information extraction Agent cooperation.
1 Introduction
@@
In the era of information abundance, constructing comprehensive knowledge graphs [36, 17, 25] has emerged as a pivotal task. The advent of LLMs, such as GPT-3 [1] and ChatGLM [4], has revolutionized natural language processing by showcasing unparalleled proficiency in understanding and generating human-like text. However, the application of these models to KGC remains an intricate challenge, as this task necessitates not only language understanding but also precise extraction of elements within the confines of predefined schemas. Recent investigations [33] reveals that the raw textual data utilized to train large language models may lack task-specific schemas, resulting in a weakened semantic grasp and structural analysis of the underlying schema. Therefore we contends that a shift from traditional parameter-based paradigms to a more nuanced approach, like Chain-of-Thought (CoT) [32, 37], can address the challenges posed by multi-step inference problems inherent in KGC. Embracing the profound insights from the Society of Mind (SOM) [16], which conceptualizes the mind as a complex system emerging from the interactions of simple components, our research explores the transformative potential of LLM-based agents in multi-agent systems. Taking inspiration from pioneering work of [13], we employ the multi-agent debate framework for collaborative self-reflection on challenging tasks. Collaboration is defined as an iterative refinement process, wherein each round generates a new answer based on prior answers and self-reflection. This iterative feedback fosters continuous improvement, making our collaborative approach adept at tackling problems that elude single-agent solutions.
@@
Specifically, our dedicated team of agents comprises experts proficient in various tasks, including named entity recognition, relation extraction, and event extraction. In our approach CooperKGC , we construct a collaborative team of agents, each specializing in distinct tasks to simulate the nuanced teamwork prevalent in human society. The integration of open interaction, expertise refinement, and adaptability to others’ opinions mirrors the foundations of a cohesive society. Our exploration into diverse collaboration strategies reveals key insights: (1) Inclusion of agents with varied expertise enhances collaboration outcomes. (2) While model hallucinations [35] may arise, effective communication among team members mitigates these drawbacks. (3) Substantial team collaboration enhances extraction results on target tasks; however, an intriguing observation emerges that increasing cooperation rounds doesn’t invariably yield superior results. In our collaborative mechanism, balancing interaction frequency ensures the expert agent’s beliefs remain undisturbed by excessive external authoritative information, aligning with fundamental theories of sociology [24, 6, 5].
2 Related work
@@
AutoKG [39] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
ChatIE [33] [Uncaptioned image] (2-agents) [Uncaptioned image] [Uncaptioned image] (2-rounds) [Uncaptioned image] [Uncaptioned image]
GPT-NER [28] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] (>3-rounds) [Uncaptioned image] [Uncaptioned image]
GPT-RE [27] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
CoT-ER [15] [Uncaptioned image] (3-agents) [Uncaptioned image] [Uncaptioned image] (3-rounds) [Uncaptioned image] [Uncaptioned image]
LM vs LM [2] [Uncaptioned image] (2-agents) [Uncaptioned image] [Uncaptioned image] (3-stages) [Uncaptioned image] [Uncaptioned image]
Multiagent Debate [3] [Uncaptioned image] (2-agents) [Uncaptioned image] [Uncaptioned image] (3-stages) [Uncaptioned image] [Uncaptioned image]
MAD [13] [Uncaptioned image] (3-agents) [Uncaptioned image] [Uncaptioned image] (3-stages) [Uncaptioned image] [Uncaptioned image]
PRD [10] [Uncaptioned image] (2-agents) [Uncaptioned image] [Uncaptioned image] (3-stages) [Uncaptioned image] [Uncaptioned image]
SPP [30] [Uncaptioned image] (>3-agents) [Uncaptioned image] [Uncaptioned image] (4-stages) [Uncaptioned image] [Uncaptioned image]
@@
Table 1: Comparison with previous methods. The upper half represents LLM-based KGC method, while the lower shows the emerging multi-agent approach.
2.1 LLM-based Knowledge Graph Construction
@@
Recent years have witnessed a surge of interest in leveraging the remarkable advancements achieved by LLMs within the realm of KG. Notably, [39] delves into the application of LLMs in KG construction and reasoning tasks. Building on this foundation, [38] integrates KG structural information into LLMs, employing self-supervised structural embedding pre-training. [33], in a novel perspective, proposes a multi-turn question and answer architecture. Furthermore, [19] pinpoints the unspecified task description as a key factor hindering the performance of contextual information extraction. To address this, a guided learning framework is introduced to enhance the extraction model’s alignment with specified guidelines. Departing from the conventional isolation of KGC as a singular task, our approach advocates a departure from such isolation by fostering collaboration among a group of expert model agents in a multi-round social environment.
2.2 Interactive Collaboration of Multiple Agents
Recent developments highlight the effectiveness of collaborative efforts among language model agents, offering potential to enhance individual LLM capabilities. Various interaction architectures have emerged, assigning agents to specific roles. For example, setups like that in [3] engage two agents in debates, enhancing factuality and reasoning albeit with increased computational costs. Similarly, [2] introduces an examiner LLM to validate claims, uncovering factual errors through division of labor. Contributions such as the ChatLLM network [7] foster dialogue and collective problem-solving among language models. Others introduce judges to summarize debates and provide conclusions [13, 34], or frameworks incorporating peer review and discussion to address self-enhancement bias [10]. Our study presents an interactive architecture tailored for a knowledge graph construction team, focusing on multiple information extractions and feedback compatibility.
Figure 1: The overview of our CooperKGC.
3 Methodology
@@
Illustrated in Figure 1, we introduces a collaborative framework CooperKGC, aimed at advancing knowledge graph construction by concurrently extracting component elements such as entities, relations, and events. Notably, our method could extend beyond the confines of the selected three tasks, offering flexibility through a dynamically formulated team collaboration network tailored to specific task requirements.
3.1 Construction of Multi-expert Agent Collaboration Network
@@
Traditional methods treat expert agents, each equipped with distinct back-ends, as isolated nodes within the collaborative network. These nodes independently contribute to task-solving through separate thinking chains, and a central adjudication node amalgamates and rectifies their responses. However, this conventional solution reveals two flaws: (1) The adjudication node, functioning as the central hub, exhibits low fault tolerance and demands substantial reasoning ability to assimilate opinions from nodes spread across diverse collaborative networks; (2) The team heavily relies on the ruling node as the sole consensus mechanism, hindering effective interactions between participants in the KGC task. In response to these limitations, we advocate a decentralized collaborative network communication scheme. Here, each expert agent backend, responsible for handling a specific task, establishes a bidirectional communication channel with any other expert agent backend. Despite the asynchronous nature of message production during practical operations, we adopt rounds as the fundamental unit of interaction to accomplish designated tasks and facilitate replica communication among expert agents. It is noteworthy that, although our approach draws inspiration from the Byzantine Fault Tolerance [8] to form a distributed network, the message records held by each agent node differ. In the process of replica communication, we implement message simplification, whereby extraction results complying with schema constraints are distilled. The formalization of the abstract collaboration network comprises three fundamental components:
@@
3.2 Customized Expert Knowledge Background.
@@
Task Definition. The task description $\mathcal{P}_{t}$ can be further decomposed into three components, as exemplified by the RE agent:(1) The first sentence of the task description, "You are an excellent expert in relation extraction." is a constant that tells the LLM that it needs to focus on the relation extraction task; (2) The second sentence defines the output format of the task: "Each result is returned as a tuple, e.g. [(head entity 1, relation type 1, tail entity 1), …]". (3) The third sentence points to a specific list of relation types : "The list of constrained relations is: [#Relation 1: [#Head Entity Type 1, #Tail Entity Type 2]…]".
@@
Algorithm 1 The Optimization Process of CooperKGC
In-context Demonstration. Some studies [27, 15, 14] show improvements in contextual learning by selecting few-shot demonstrations based on similarity. Our contextual prompts $\mathcal{P}_{c}$ are introduced as N-way K-shot sampling of the demonstration samples $\mathcal{M}=N\times K$, providing direct evidence about the task and references to predictions. However, limited by the input tokens of the LLMs, a single prompt may not contain all supported instances, so we use a sentence embedding similarity-based approach to select the $\mathcal{M}$ examples with the closest Euclidean distance as contexts.
3.3 Collaborative Interaction Optimization
In the context of team collaboration optimization, the need for meticulous decomposition design diminishes, thus we reach to the periphery of the age-old adage, "Two heads are better than one." As shown in Algorithm 1, after collecting replicas by other expert agents, we further provide collaboration prompts $\mathcal{P}_{v}$: The relation extraction answer you gave in the last round of collaboration was "##LAST_ROUND_RESULT##". The answer given by the NER expert agent was "##NER_RESULT##", The EE expert agent was "##EE_RESULT##". You should refer to other members to revise your answer."
4 Experiments
  • RQ3: How effective is the proposed CooperKGC in extracting different types of entities, relations and events?
4.1 Experiment Settings
4.1.1 Dataset.
As to the NER task, we conduct experiments on the following popular benchmark: Conllpp [31], OntoNotes5.0 [20] and MSRA [9]. For RE task, we conduct experiments on the following popular benchmark: NYT11-HRL [23], Re-TACRED [21], and DuIE2.0 [11]. For EE task, there are two standard datasets: ACE05 [26] and DuEE1.0 [12].
4.1.2 Baselines.
In our experimental framework, we opt for AutoKG [39] as the implementation of Vanilla LLMs for KGC realm, which defines an end-to-end extraction workflow through the manual templates. Expanding on this foundation, ChatIE [33] refines the extraction process using a two-round method. Taking RE as an example, this method entails the initial extraction of the relation, followed by the output of the associated entity span. This sequential approach mirrors a cognitive model’s thought process, explicitly delineating the steps of task decomposition. Further, CoT-ER [15] introduces an explicit evidence reasoning method, characterized by three rounds of processing. In the first and second rounds, the LLM is required to output concept-level entities corresponding to head and tail entities. Subsequently, in the third round, the extraction of relevant entity spans occurs, establishing a specific relationship between these two entities with explicit evidence.
4.2 Performance Comparison with SOTA (RQ1)
Table 2: F1-score results for 3 KGC tasks (NER, RE, EE) on the 8 datasets.